if 18c2r=153*13c2r-5 what is the value of r?
Answer by tutorcecilia(2144)

18c^r=153(13c^r-5)
18c^2=1989c^2r-765
18c^2+765=1989c^2-765+765
18c^2+765=1989c^2
18c^2-18c^2+765=1989c^2-18c^2
765=1989c^2-18c^2
765=1971c^2r
765/1971c^2=1971c^2r/1971c^2
765/1971c^2=r
85/219c^2=r
.
Check by plugging (85/219c^2=r) back into the original equation.